OpenStudy (anonymous):

find dy/dx from y = 2x^3 * ln(ax)

zepdrix (zepdrix):

yayy another product ruleeee

zepdrix (zepdrix):

What part you getting stuck on? :O

OpenStudy (anonymous):

\[(fg)'=f'g+fg'\]

OpenStudy (anonymous):

with \[f(x)=2x^3,f'(x)=6x^2,g(x)=\ln(ax), g'(x)=\frac{1}{x}\]

OpenStudy (anonymous):

@ zepdrix, i just need to check my answers lol... i keep thinking that i did things wrong

zepdrix (zepdrix):

Oh i see :)

OpenStudy (anonymous):

isn't it chain rule with ln(ax)?

OpenStudy (anonymous):

i meant product rule?

zepdrix (zepdrix):

chain rule? :o oh yes i suppose it is. \[[\ln(ax)]'=\frac{ 1 }{ ax }(ax)'=\frac{ 1 }{ ax }a=\frac{ 1 }{ x }\] But you still get what satalite posted :3 after simplifying.

zepdrix (zepdrix):

Well hopefully they intended for A to be a constant, which is why we wouldn't apply the product rule :O it's not a variable thingy

OpenStudy (anonymous):

well there's no where that says a is a constant :(